OpenStudy (anonymous):

Simplify: √a² and √36a⁴b¹⁰

OpenStudy (anonymous):

\[\sqrt{a^2}=a\]

OpenStudy (anonymous):

who did you get theses?

OpenStudy (anonymous):

just a sec...

OpenStudy (anonymous):

ok thankss

OpenStudy (anonymous):

soz its\[6 \times a^2\times b^5\]

OpenStudy (anonymous):

because root 36 =6 \[(a^4)^{1/2} = a^2 \] \[(b^{10})^{1/2} = b^5\]

OpenStudy (anonymous):

using the laws of powers the sqrt of 8 is \[8^{1/2}\] do you see

OpenStudy (anonymous):

yesss. thank you!

OpenStudy (anonymous):

:D I love that yessss moment